Circular No. 6631

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) G. Cremonese, Astronomical Observatory, Padua, on behalf of the Eu-

ropean Hale-Bopp Team, reports the detection of an extensive neutral Na tail structure associated with this comet: “The structure is visible on images obtained by D. Pollacco, Isaac Newton Group, with the wide-field CoCAM instrument (aperture 50 mm, 2k × 1k EEV CCD, 26′′/pixel) at the Observatorio del Roque de los Muchachos, La Palma. The following description is based on two 10-min exposures obtained through an Na D filter on Apr. 16.88 UT under mediocre sky conditions. The filter (λ_{0} = 589.2 nm, FWHM = 2.1 nm) has no leak at the 0.02-percent level over the entire sensitivity range (300–1100 nm) of the CCD. The images display two main tails. The first is a very linear feature 6^{o}.6 long and ≤ 10′ wide with parallel edges over its entire length and a very sharp south edge at p.a. 52^{o} (3^{o} north of the projected antisolar direction). The second is a broad diffuse northern tail extending to the edge of the frame, i.e., to > 6^{o}.5. While the second tail coincides with the dust tail as seen in an adjacent continuum filter, the first tail has no such counterpart, and we identify it as a neutral Na tail. In particular, near-simultaneous imaging with an H_{2}O^{+} filter (λ_{0} = 619.3 nm, FWHM = 4.5 nm) displays an ion tail of entirely different and more complex morphology with a main ray at p.a. 46^{o}. We believe that this is the first clear separation of a neutral tail of a comet from the dust and ion tails. The existence of this third type of cometary tail was confirmed in additional CoCAM images obtained on Apr. 17.8 UT.”

GRB 970228 M. R. Metzger, J. L. Cohen, J. P. Blakeslee, S. R. Kulkarni, S. G. Djor-

govski and C. C. Steidel, California Institute of Technology; and D. A. Frail, National Radio Astronomy Observatory, report: “Moderately deep images of the region near the optical variable reported on IAUC 6584 were obtained on Apr. 5.24 and 6.27 UT with LRIS on the Keck II 10-m telescope. The faint source noted by Groot et al. and by Metzger et al. on IAUC 6588 and observed later by HST (Sahu et al., IAUC 6606, 6619) was measured in a combined image (15 min total exposure time) at R = 24.9 ± 0.3. In the combined image (with seeing 0′′.7 FWHM), the source does not appear significantly extended. Direct comparison with the Mar. 6.32 image (Metzger et al., IAUC 6588) indicates that the total R magnitude has faded by 1.0 ± 0.4 mag and that the component causing the earlier images to appear extended has faded below detection level.”

1997 April 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden